OpenStudy (xochital):

Part 1: Solve the inequality -3x + 4 x + 8 and show all of your work. You can use <= to represent the less than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1.

OpenStudy (xochital):

-3x + 4 <= x + 8

OpenStudy (anonymous):

- 3x + 4 <= x + 8 - 3x - x <= 8 - 4 -4x <= 4 -4x/4 >= 4/-4 (switch the sign when dividing by a negative) x >=-1 graphing....on a number line... closed circle on -1 shaded to the right

OpenStudy (xochital):

thank you! i need help with one more... Part 1: Use complete sentences to describe a real-world scenario that could be represented by the inequality 5x + 2y >= 45. Part 2: Choose one ordered pair that is a solution to the given inequality and explain what that ordered pair means in the context of your real-world scenario.

OpenStudy (anonymous):

I am not so good at graphing with an x and a y.....sorry

OpenStudy (xochital):

that's fine, thanks for your help :)

OpenStudy (anonymous):

sorry that I couldn't help more :(

OpenStudy (xochital):

don't worry about it! :)

OpenStudy (anonymous):

thx for the medal :)